# BMGQ: A Bottom-up Method for Generating Complex Multi-hop Reasoning Questions from Semi-structured Data

### v1

@@ Abstract
To address this, we present an automated framework for generating high-difficulty, training-ready multi-hop questions from semi-structured knowledge sources. The system (i) grows diverse, logically labeled evidence clusters through Natural Language Inference (NLI)-based relation typing and diversity-aware expansion; (ii) applies reverse question construction to compose oblique cues so that isolated signals are underinformative but their combination uniquely identifies the target entity; and (iii) enforces quality with a two-step evaluation pipeline that combines multi-model consensus filtering with structured constraint decomposition and evidence-based matching. The result is a scalable process that yields complex, retrieval-resistant yet verifiable questions suitable for SFT/RL training as well as challenging evaluation—substantially reducing human curation effort while preserving the difficulty profile of strong evaluation benchmarks.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress across many natural language processing tasks, yet their ability to conduct deep research involving multi-hop retrieval and reasoning remains limited. This capability is crucial for solving problems where answers cannot be found through a single lookup but require integrating multiple pieces of indirect, cross-domain evidence. While existing multi-hop QA training datasets provide useful testbeds, most of them rely on relatively shallow reasoning chains, offering limited value for training models that can handle genuinely complex retrieval and reasoning tasks.
@@
Motivated by these gaps, we introduce an automated framework for constructing high-difficulty multi-hop datasets suitable for both training and evaluation. Our approach transforms semi-structured knowledge sources into structured evidence clusters, builds diverse and logically coherent evidence clusters, and employs a reverse construction strategy that composes indirect clues to ensure that the answer is uniquely identifiable only when all clues are considered together. Finally, a dedicated quality evaluation system combines multi-model consensus filtering with structured constraint decomposition and evidence-based matching, ensuring that each retained question is both challenging and verifiably correct.
By combining the difficulty profile of high-end evaluation benchmarks with the scalability of automated synthesis, our framework enables the creation of large, training-ready multi-hop datasets that are capable of truly stressing and improving the deep retrieval and reasoning abilities of large language models.
@@ 2 Related Works
2.1 Early Benchmarks: Progress and Limitations
A series of early multi-hop QA benchmarks laid the foundation for research on multi-hop reasoning. Among them, several datasets progressively expanded the scope of reasoning—from basic two-hop evidence combination to more compositional chains and structured knowledge integration—marking an important trajectory in the evolution of the field. However, despite these advances, they remain fundamentally limited in reasoning depth, retrieval difficulty, and answer uniqueness, making them insufficient for training or evaluating models designed for complex, open-domain reasoning.
HotpotQA [1] represents a seminal milestone, introducing large-scale multi-document QA with annotated supporting sentences to encourage explicit reasoning. While it successfully established the importance of evidence combination, most of its questions involve only two reasoning hops and can often be solved through shallow retrieval heuristics or surface-level lexical matching, rather than genuine multi-step inference.
Building on this foundation, ComplexWebQuestions [2] sought to increase reasoning complexity by automatically generating compositional questions from knowledge base paths. This approach broadened the scope beyond simple paragraph linking and introduced more structured multi-hop reasoning. However, its template-based construction limited linguistic variability, and the early versions suffered from train-test leakage, ultimately constraining its diagnostic power.
MuSiQue [3] further advanced this line by deliberately composing single-hop facts into multi-hop chains and introducing mechanisms to reduce spurious shortcuts. Compared to its predecessors, it provides richer compositional structures and greater control over reasoning paths. Yet, most questions still require relatively short reasoning chains, and the dataset does not explicitly enforce retrieval difficulty or the “hard-to-search but easy-to-verify” property crucial for deep research scenarios.
Other datasets, such as QASC [4] and 2WikiMultiHopQA [5], extend this trend into specific domains or structured-unstructured hybrids. QASC focuses on elementary science reasoning through sentence combination, while 2Wiki combines Wikidata triples with Wikipedia passages to provide explicit reasoning path annotations. Though each brings incremental improvements, their domain restrictions, template-driven construction, and reliance on structured or encyclopedic sources limit their ability to model realistic open-domain retrieval settings.
Overall, while these benchmarks were instrumental in shaping the multi-hop QA landscape, they largely emphasize shallow reasoning, short-hop paths, and constrained domains. As a result, they fall short of capturing the full complexity, ambiguity, and uniqueness constraints that characterize real deep-research multi-hop retrieval and reasoning tasks.
2.2 BrowseComp: A Shift Toward “Hard-to-Search, Easy-to-Verify”
The BrowseComp dataset [6] represents a significant step forward toward realistic multi-hop evaluation. Unlike earlier benchmarks, BrowseComp is designed around the principle of “hard to search, easy to verify.” Individually, these clues are too vague to retrieve the answer directly, but only by jointly reasoning over multiple pieces of information can the search space be narrowed to a single correct answer.
BrowseComp questions exhibit several distinctive characteristics:
  • Deep and wide reasoning chains: Solving them often requires traversing multiple knowledge domains and integrating scattered evidence across several web pages.
  • Fuzzy and indirect clues: Signals are deliberately vague or partial, demanding reasoning over context rather than direct lookup.
  • Hard-to-search but easy-to-verify: While the answer is difficult to retrieve directly, once found, it is short, concrete, and objectively verifiable.
  • Bias toward niche domains: Many answers are rare entities or domain-specific facts, discouraging memorization and forcing deeper retrieval strategies.
Building on this idea, BrowseComp-ZH [7] extends the framework to the Chinese web domain, aiming to benchmark multilingual web-browsing capabilities of large language models. However, our internal manual evaluation of both BrowseComp and BrowseComp-ZH reveals that a substantial portion of their questions lack strict answer uniqueness or contain ambiguities in clue interpretation. In many cases, multiple plausible answers can satisfy the same set of fuzzy constraints, highlighting the difficulty of consistently encoding high reasoning complexity while ensuring unambiguous ground truths.
This observation underscores a crucial gap in the current landscape of multi-hop datasets. Both BrowseComp and BrowseComp-ZH were intentionally designed as evaluation benchmarks, not training corpora, aiming to push the limits of retrieval and reasoning capabilities of large language models. While they effectively capture realistic reasoning difficulty, our manual inspection further reveals that many questions exhibit imperfect answer uniqueness, making them less suitable for direct use in training scenarios where strict supervision is required. This highlights the need for automated construction pipelines with explicit quality control that can generate large-scale datasets matching the difficulty of BrowseComp while ensuring unambiguous and verifiable ground truths for model training.
2.3 Automatic Construction of Multi-Hop Datasets
As the demand for large-scale, high-quality multi-hop datasets grows, researchers have explored various automatic construction frameworks to alleviate the prohibitive costs of manual curation. These approaches aim to generate questions and reasoning chains directly from structured knowledge bases, unstructured text, or a combination of both. However, existing methods remain limited in scope and fall short of replicating the depth, ambiguity, and reasoning complexity seen in challenging datasets such as BrowseComp[6].
Template- and Knowledge-Graph-Based Methods.
Early efforts in automatic dataset construction relied heavily on rule-based templates and knowledge graph traversal. For instance, ComplexWebQuestions [2] generated compositional questions by expanding seed questions in WebQuestionsSP with additional relations from Freebase. This approach enabled scalable generation but produced synthetic and predictable question structures, with reasoning typically confined to simple relation chaining. Similarly, 2WikiMultiHopQA [5] constructed multi-hop questions by linking Wikidata triples and retrieving supporting passages from Wikipedia, but its reliance on predefined templates and explicit relation paths limited linguistic diversity and often failed to capture the subtle reasoning required in real-world retrieval. While these methods demonstrated the feasibility of automatic construction, they struggled with several key challenges:
  • Limited reasoning diversity: Most generated questions followed fixed reasoning patterns (e.g., bridge or comparison), leading to repetitive question structures.
  • Lack of ambiguity and indirectness: Template generation typically results in explicit, easily searchable questions, lacking the fuzzy clues and oblique phrasing that characterize harder multi-hop tasks.
  • Weak retrieval difficulty: Because many questions can be answered by direct lookup of a single knowledge base triple, they fail to replicate the retrieval complexity of open-domain scenarios.
HopWeaver: Toward Automated Multi-Hop Question Generation.
A more recent and influential attempt at automatic dataset construction is HopWeaver [8], which introduces a retrieval-augmented generation framework to synthesize multi-hop questions. HopWeaver systematically constructs bridge-type and comparison-type questions by identifying semantically related entity pairs and retrieving intermediate evidence passages from Wikipedia. It then uses large language models to compose natural-language questions that integrate information across multiple hops. This work represents a substantial advance beyond purely template-based approaches, as it leverages both retrieval signals and generative models to produce more natural and varied questions. However, despite these improvements, HopWeaver still exhibits several critical limitations:
  • Restricted reasoning types: The framework focuses primarily on bridge and comparison questions, leaving out many reasoning forms involving casual, compositional, or conditional relations.
  • Shallow reasoning depth: Most generated questions remain within two to three hops, lacking the depth and cross-domain traversal seen in BrowseComp-style problems.
  • Insufficient ambiguity: HopWeaver does not explicitly enforce fuzzy clue construction or the ”hard-to-search but easy-to-verify” principle, resulting in questions that are still vulnerable to direct retrieval.
  • Limited control over answer uniqueness: The framework does not guarantee that a question points uniquely to a single answer without external disambiguation, which is critical for training and evaluating reasoning models.
In summary, existing automatic construction frameworks — from template-based systems to retrieval-augmented generation like HopWeaver — represent important milestones in scaling multi-hop dataset creation. Yet they remain insufficient for producing BrowseComp-level questions that combine deep reasoning chains, ambiguous and indirect clues, domain-crossing knowledge transitions, and strict answer uniqueness. Bridging this gap requires a fundamentally new approach that integrates structured evidence graph construction, bottom-up question synthesis, and rigorous quality evaluation — capabilities that motivate the framework proposed in this work.
2.4 Dataset Quality Control in Automatic Generation
While automatic generation greatly reduces the cost of dataset construction, it also introduces significant challenges in ensuring the correctness, uniqueness, and reasoning validity of the resulting question–answer pairs. Poorly controlled generation often leads to issues such as ambiguous questions, incorrect answers, shallow reasoning chains, or spurious shortcuts that undermine the value of the dataset. As a result, effective post-generation quality control is essential for building reliable multi-hop datasets.
Rule-Based and Discriminative Filtering Approaches.
Early approaches to quality assurance primarily relied on rule-based heuristics or discriminative models applied after question generation. For instance, ComplexWebQuestions employed rule-based checks to discard syntactically malformed questions or those with mismatched entities. Other systems used lightweight classifiers to verify whether retrieved documents supported the predicted answers.
LLM-Based Evaluation and Self-Verification.
Recent work has explored leveraging large language models themselves as evaluators, introducing more semantic and reasoning-aware quality control mechanisms. SelfCheckGPT [9] uses a model to assess the factual consistency of its own generations by generating multiple paraphrases and measuring agreement. Similarly, LLM-as-a-Judge approaches [10] use language models as evaluators to assess factuality, reasoning soundness, or answer correctness without explicit labels. These methods represent a step forward in capturing semantic errors and deeper reasoning flaws beyond surface matching.
However, most existing LLM-based evaluation frameworks are model-centric and rely on a single model’s judgment, which introduces bias and lacks robustness. They often do not incorporate structured representations of the question’s internal logic, nor do they enforce explicit constraints on time, location, or entity properties. As a result, they cannot fully guarantee that a question is logically coherent, uniquely solvable, and well-aligned with available evidence.
To address these limitations, we propose a structured Data Quality Evaluation System that integrates multi-model consensus with constraint-based verification. Rather than treating quality control as a single-step classification, our approach first filters ambiguous questions through model agreement and then verifies candidate answers via decomposed, evidence-grounded predicates, ensuring logical soundness and answer uniqueness.
Therefore, against this backdrop, our work contributes a novel pipeline for automatic multi-hop dataset construction that explicitly addresses these gaps. We design a methodology capable of producing BrowseComp-level questions at scale for training purposes, and we augment it with a Data Quality Evaluation System that combines multi-model consensus, structured clue decomposition, and evidence-based verification. This integration ensures that the resulting dataset is not only large and diverse, but also auditable, logically coherent, and uniquely solvable, laying the groundwork for advancing both training and evaluation of reasoning-capable large language models.
3 Methodology
@@
Our methodology for constructing a multi-hop reasoning dataset follows a structured four-stage pipeline, as illustrated in Figure 1. In Part 1 (Data Sources & Adaptation), raw Wikipedia and Wikidata data are adapted into a lightweight relational database and pre-constructed evidence network. In Part 2 (Node Information Construction), we retrieve raw page content, perform text preprocessing, remove non-entity evidence, and link candidate entities with their supporting paragraphs. In Part 3 (Evidence Chain Construction), relation classification guarantees the validity of edges, diversity-aware evaluation mitigates semantic redundancy, and breadth-first expansion yields a balanced and interpretable evidence cluster. Finally, in Part 4 (Question Construction & Optimization), a bottom-up reverse generation strategy constructs the initial question, followed by multi-round optimization to enhance abstraction, difficulty, and uniqueness. Together, these stages enable the automatic generation of high-quality, logically coherent, and challenging multi-hop questions grounded in real-world knowledge.
@@ 3.2 Node Information Construction
Once the structured database has been established, the next step is to construct an informative contextual environment for a given seed entity. This process aims to identify high-quality candidate entities and their corresponding evidence passages, which serve as the foundation for multi-hop evidence chain construction.
The key challenge arises from the fact that a single Wikipedia page may contain hundreds or even thousands of outbound links, many of which correspond to generic terms, abstract concepts, or weakly related references. Blindly expanding all such links leads to semantic drift [11][12] — a phenomenon where the reasoning path gradually loses topical or contextual relevance, resulting in misleading or meaningless connections. To mitigate this issue, we introduce a multi-stage pipeline that integrates preprocessing, entity recognition, and evidence alignment. An annotated example illustrating the prevalence of irrelevant or weakly connected links in raw Wikipedia pages is provided in Appendix A.

### v2

@@ Abstract
To address this, we present BMGQ, a bottom-up automated method for generating high-difficulty, training-ready multi-hop questions from semi-structured knowledge sources. The BMGQ system (i) grows diverse, logically labeled evidence clusters through Natural Language Inference (NLI)-based relation typing and diversity-aware expansion; (ii) applies reverse question construction to compose oblique cues so that isolated signals are underinformative but their combination uniquely identifies the target entity; and (iii) enforces quality with a two-step evaluation pipeline that combines multi-model consensus filtering with structured constraint decomposition and evidence-based matching. The result is a scalable process that yields complex, retrieval-resistant yet verifiable questions suitable for SFT/RL training as well as challenging evaluation—substantially reducing human curation effort while preserving the difficulty profile of strong evaluation benchmarks.
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable progress across many natural language processing tasks, yet their ability to conduct deep research involving multi-hop retrieval and reasoning remains limited. This capability is crucial for solving problems where answers cannot be found through a single lookup but require integrating multiple pieces of indirect, cross-domain evidence. While existing multi-hop QA training datasets provide useful testbeds, most of them[1][2] rely on relatively shallow reasoning chains, offering limited value for training models that can handle genuinely complex retrieval and reasoning tasks.
@@
Motivated by these gaps, we introduce BMGQ, a bottom-up automated framework for constructing high-difficulty multi-hop datasets suitable for both training and evaluation. BMGQ transforms semi-structured knowledge sources into structured evidence clusters, builds diverse and logically coherent evidence clusters, and employs a reverse construction strategy that composes indirect clues to ensure that the answer is uniquely identifiable only when all clues are considered together. Finally, a dedicated quality evaluation system combines multi-model consensus filtering with structured constraint decomposition and evidence-based matching, ensuring that each retained question is both challenging and verifiably correct.
By combining the difficulty profile of high-end evaluation benchmarks with the scalability of automated synthesis, BMGQ enables the creation of large, training-ready multi-hop datasets that are capable of truly stressing and improving the deep retrieval and reasoning abilities of large language models. To support reproducibility, we release a partial subset of the BMGQ dataset, including only the final questions and their seed answers, at https://huggingface.co/datasets/Fayer/BMGQ-MultiHop-Sample
@@ 2 Related Works
2.1 Constructing Deep, Ambiguous, and Uniquely Solvable Multi-Hop Questions
We study the problem of automatically constructing large-scale multi-hop QA datasets that match the reasoning difficulty of human-curated benchmarks such as BrowseComp while ensuring strict answer uniqueness. Unlike traditional QA tasks, the goal is not merely to answer a question, but to generate a question that:
  1. 1.
    Requires deep multi-step reasoning across a chain of heterogeneous evidence.
  2. 2.
    Contains ambiguous or indirect clues that prevent shallow retrieval or direct lookup.
  3. 3.
    Remains uniquely solvable, meaning only one entity in the evidence space satisfies all constraints.
  4. 4.
    Is verifiable against structured evidence, allowing downstream models to check correctness.
We formalize the dataset construction problem as generating a natural-language question $Q$ from a structured evidence graph $G=(V,E)$ induced by a seed entity $s$ – which also serves as the answer $A$ to be recovered. Each question imposes a set of constraints $C(Q)$ over nodes and attributes in $G$. The target answer $A$ must satisfy all constraints:
$A=\{\,v\in V\mid v\models C(Q)\,\}.$
To ensure strict answer uniqueness, the question must satisfy:
$|A|=1.$
At the same time, the reasoning path required to reach $A$ must be non-trivial:
$\text{depth}(C(Q))\geq n,$
with $n$ representing a configurable minimal clue depth.
This formulation highlights why automatic construction is difficult: deep multi-hop reasoning, clue ambiguity, cross-domain transitions, and strict uniqueness must be satisfied simultaneously, yet existing datasets or generation frameworks do not jointly enforce these constraints.
2.2 Existing Benchmarks for Multi-Hop Reasoning
Building directly upon the problem defined in Section 2.1, we evaluate how existing multi-hop QA benchmarks relate to the goal of constructing deep, ambiguous, and uniquely solvable questions. Although these datasets have significantly advanced multi-hop reasoning research, none of them are designed to satisfy all required properties—deep reasoning depth, indirect clue construction, retrieval difficulty, and strict answer uniqueness—simultaneously.
Early benchmarks such as HotpotQA[1], ComplexWebQuestions[2], MuSiQue[3], QASC[4], and 2WikiMultiHopQA [5] laid the groundwork for multi-hop QA. HotpotQA introduced multi-document reasoning with annotated supporting sentences, but its reasoning chains are typically shallow and often solvable via lexical matching rather than true multi-step inference. ComplexWebQuestions expanded compositional complexity through knowledge-graph path generation, yet its template-based design limits linguistic diversity and yields predictable reasoning structures. MuSiQue aimed to mitigate shortcut exploitation by composing single-hop facts into longer chains, but the majority of its questions remain short and lack enforced retrieval difficulty. Domain-focused and hybrid-structured datasets such as QASC and 2WikiMultiHopQA broadened coverage but still depend heavily on explicit facts or templates. Collectively, these benchmarks remain limited by shallow reasoning, explicit clues, and a lack of guarantee on answer uniqueness.
A more recent and significantly harder benchmark, BrowseComp[6], represents a distinct shift in design philosophy. BrowseComp constructs questions that are deliberately vague when examined clue-by-clue, forcing deep multi-hop retrieval across multiple domains, yet the final answer is short and objectively verifiable. This “hard-to-search, easy-to-verify” paradigm better captures the challenges of real-world deep research tasks, requiring models to integrate scattered, indirect signals rather than relying on direct lookup. However, despite its conceptual advances, our internal manual evaluation suggests that many BrowseComp questions still exhibit imperfections in uniqueness: multiple plausible answers may satisfy the same fuzzy constraints. Moreover, BrowseComp is intentionally released as an evaluation benchmark, not a scalable training corpus, making it unsuitable for supervised finetuning or reinforcement learning.
In summary, while existing multi-hop datasets collectively push the field toward more realistic reasoning, they fall short of the requirements outlined in Section 2.1. Early benchmarks lack depth, ambiguity, and retrieval difficulty, whereas BrowseComp achieves substantial difficulty but still does not guarantee strict answer uniqueness or provide a method for scalable dataset construction. This persistent gap highlights the need for a principled construction framework capable of producing deep-reasoning, ambiguous-yet-uniquely-solvable questions at scale—an objective that motivates the methodology presented in this work.
2.3 Automatic Multi-Hop Dataset Construction Frameworks
Given the limitations of existing multi-hop benchmarks, researchers have increasingly explored automatic question construction as a scalable alternative to manual curation. Early efforts primarily relied on rule-based templates or knowledge graph traversal, aiming to generate multi-hop questions by expanding relations in structured datasets. For example, ComplexWebQuestions constructs compositional questions by extending WebQuestionsSP using Freebase relation paths. Although this approach enables large-scale synthesis, its dependence on fixed templates results in predictable linguistic forms and shallow, deterministic reasoning chains. Similar KG-driven pipelines—such as those used in 2WikiMultiHopQA – combine Wikidata triples with supporting Wikipedia passages, but remain constrained by explicit relation paths that lack the ambiguity and retrieval difficulty required in deep multi-hop contexts.
To move beyond template rigidity, several works have explored retrieval-augmented generation. The most influential of these is HopWeaver [7], which synthesizes bridge and comparison questions by identifying semantically connected entity pairs, retrieving intermediate passages from Wikipedia, and prompting an LLM to compose a question integrating those hops. HopWeaver demonstrates that automatic multi-hop generation can leverage textual evidence and large language models to produce more natural questions than purely symbolic pipelines. However, its reasoning patterns remain limited to a narrow set of hop types (mainly bridge and comparison), and most generated questions only require two to three hops. Moreover, HopWeaver does not explicitly enforce controlled ambiguity or guarantee strict answer uniqueness—properties that are central to the construction of hard-to-search yet uniquely solvable questions. As a result, although HopWeaver marks an important step forward, it still falls short of producing BrowseComp-level difficulty in a scalable and principled way.
Taken together, existing automatic construction methods provide valuable insights but do not jointly address the core challenges required for generating deep multi-hop reasoning datasets. They either rely on shallow templates, produce limited reasoning diversity, or lack robust verification pipelines to ensure correctness and uniqueness. These gaps motivate the development of a new construction framework—one that leverages structured evidence graphs, bottom-up reverse synthesis, and explicit quality evaluation—to generate BrowseComp-level questions at scale.
2.4 Remaining Gaps and Our Contribution
From the above analysis, we identify a critical gap: there exists no automatic pipeline capable of generating BrowseComp-level hard multi-hop questions with explicit, auditable evidence and guaranteed answer uniqueness. To bridge this gap, we propose a new automated construction framework that integrates:
  • structured evidence graph building,
  • depth-controlled clue selection,
  • bottom-up question synthesis,
  • graph-based structural sanity checks, and
  • a two-layer Data Quality Evaluation System combining (i) model-based coarse filtering and (ii) predicate-level constraint verification.
Formally, the goal of our system is:
$\text{Construct}(Q)\quad\text{s.t.}\quad\begin{cases}\text{depth}(C(Q))\geq n,\\ \text{ambiguity}(Q)\geq\tau_{1},\\ \text{retrieval\text{-}difficulty}(Q)\geq\tau_{2},\\ \bigl|\{\,v\in V\mid v\models C(Q)\,\}\bigr|=1.\end{cases}$
This framework unifies graph construction, cue abstraction, question synthesis, and evidence-based verification. It enables the construction of BrowseComp-grade multi-hop reasoning datasets at scale, while ensuring strict answer uniqueness, logical coherence, and structured auditability.
3 BMGQ Framework: Bottom-up Method for Generating Complex Multi-hop Reasoning Questions
@@
Our methodology for constructing a multi-hop reasoning dataset follows a structured four-stage pipeline, collectively referred to as BMGQ (Bottom-up Method for Generating Complex Multi-hop Reasoning Questions), as illustrated in Figure 1. In Part 1 (Data Sources & Adaptation), semi-structured data are adapted into a lightweight relational database and pre-constructed evidence network. In Part 2 (Node Information Construction), we retrieve raw page content, perform text preprocessing, remove non-entity evidence, and link candidate entities with their supporting paragraphs. In Part 3 (Evidence Chain Construction), relation classification guarantees the validity of edges, diversity-aware evaluation mitigates semantic redundancy, and breadth-first expansion yields a balanced and interpretable evidence cluster. Finally, in Part 4 (Question Construction & Optimization), a bottom-up reverse generation strategy constructs the initial question, followed by multi-round optimization to enhance abstraction, difficulty, and uniqueness. Together, these components form the core of BMGQ, enabling the automatic generation of high-quality, logically coherent, and challenging multi-hop questions grounded in real-world knowledge.
@@ 3.2 Node Information Construction
Once the structured database has been established, the next step is to construct an informative contextual environment for a given seed entity [^1] The target answer of the constructed question. . This process aims to identify high-quality candidate entities and their corresponding evidence passages, which serve as the foundation for multi-hop evidence chain construction.
The key challenge arises from the fact that a single webpage may contain hundreds or even thousands of outbound links, many of which correspond to generic terms, abstract concepts, or weakly related references. Blindly expanding all such links leads to semantic drift [8][9] — a phenomenon where the reasoning path gradually loses topical or contextual relevance, resulting in misleading or meaningless connections. To mitigate this issue, we introduce a multi-stage pipeline that integrates preprocessing, entity recognition, and evidence alignment. An annotated example illustrating the prevalence of irrelevant or weakly connected links in raw Wikipedia pages is provided in Appendix A.